OpenStudy (bluestar70556):

If f(x) = (5x)/(x + 3), find f(x)^-1. Please help, I am stumped, I have a bit of it solved but I'm stuck at one part: If we make f(x) = y, switch y with x and then solve for y we get: x = (5y)/(y+3) x(y + 3) = 5y It's here I get stumped, because I can bring over the y and get (xy + 3x)/(y) = 5, but where do I go from here?

Nnesha (nnesha):

good. \[\huge\rm \frac{xy+3x}{y}=5\] is same as \[\huge\rm \frac{xy}{y} +\frac{3x}{y}=5\] rewrite it as a separate fraction

Nnesha (nnesha):

\[\large\rm \frac{ x\cancel{y} }{\cancel{y} } +\frac{3x}{y}=5\] \[x+\frac{3x}{y}=5\] solve for y

OpenStudy (bluestar70556):

Okay so now I'm at \[\frac{ 3x }{ y }=5-x\] How do I bring over the 3x? I don't think I can just multiply... Or can I?

Nnesha (nnesha):

well how would you get rid of the fraction ?

OpenStudy (bluestar70556):

Hang on I think I got it \[\left(\begin{matrix}1 \\ 3x\end{matrix}\right)\left(\begin{matrix}3x \\ y\end{matrix}\right) = 5-x\] Which then would give me \[y = 5-x \left(\begin{matrix}1 \\ 3x\end{matrix}\right)\] Which simplifies to \[y = \frac{ 5-x }{ 3x}\] Is that correct?

Nnesha (nnesha):

\[\frac{ 1 }{ \cancel{3x} } \cdot \frac{\cancel{3x}}{y} =\frac{1}{y} \] \[\frac{1}{y} \cancel{= }y\]

Nnesha (nnesha):

y/1= y 1/y is just 1/y we can move y to the right side.. right? \[\huge\rm \frac{ 3x }{ y }=5-x\] \[\huge\rm y \cdot \frac{ 3x }{ y }= (5-x) \cdot y\] \[\huge\rm \cancel{y} \cdot \frac{ 3x }{\cancel{ y} }= (5-x) \cdot y\] \[\rm 3x= \color{Red}{y}(5-x)\]

OpenStudy (bluestar70556):

OH Okay, I see now. So then we can bring over 5 - x instead... So that gives us \[\frac{ 3x }{ 5-x } = y\] And then we flip everything back to their original variables meaning: \[f(x)^-1 = \frac{ 3x }{ 5-x }\] And if I graph that it is in fact the inverse, fantastic, thank you so so much!

Nnesha (nnesha):

yw o^_^o